OpenStudy (anonymous):

What transformations change the graph of (f)x to the graph of g(x)? f(x) = x² ; g(x) = (x + 7)² + 9

OpenStudy (anonymous):

The graph of g(x) is the graph of f(x) translated to the down 7 units and right 9 units. The graph of g(x) is the graph of f(x) translated to the up 7 units and left 9 units. The graph of g(x) is the graph of f(x) translated to the right 7 units and down 9 units. The graph of g(x) is the graph of f(x) translated to the left 7 units and up 9 units. Please explain!

OpenStudy (shamil98):

D.

OpenStudy (azureilai):

the "-7" within the parentheses means that it is translating right 7. The "+9" at the end always means transltaion up or down, and since it is positive, it means 9 up.

OpenStudy (shamil98):

(x+7)^2 = left 7 units, . + 9 is up 9 units.

OpenStudy (shamil98):

(x+7)^2 is a phase shift.

OpenStudy (anonymous):

if we add/subtract from domain then its shift to left/right and if from co domain then up and down so answer is D

OpenStudy (anonymous):

left 7 and up 9 so it would be D since the equation for vertex form is y=(x-h)^2+k (h,k) being the vertex since the parent function has a vertex of (0,0) it just becomes f(x)=x^2 so in order for the number inside (x-h)^2 to become positive, h must be negative therefore shifting left and since k is positive it would shift up

OpenStudy (anonymous):

memorize your basic functions :)